11. For the reaction: D ! products, the Rate Law is: Rate = k [D]2.
If the concentration of D falls from 0.890 M to 0.560 M in 12 minutes,
what is the half life of the reaction?
Answer #1

we have:

[D]o = 0.89 M

[D] = 0.56 M

t = 12.0 min

use integrated rate law for 2nd order reaction

1/[D] = 1/[D]o + k*t

1/(0.56) = 1/(0.89) + k*12

1.786 = 1.124 +k*12

k*12 = 0.6621

k = 5.518*10^-2 M-1.min-1

Given:

k = 5.518*10^-2 M-1.min-1

use relation between rate constant and half life of 2nd order reaction

t1/2 = 1/([D]o*k)

= 1/(0.89*5.518*10^-2)

= 20.36 min

Answer: 20.4 min